What is the value of Sin inverse 2 root 2 upon 3? Plese tell

Grade:12

1 Answers

Snehal Jadhav
5 years ago
It is 75(degrees) or 5\pi/12.
 
As, the value of sin(5\pi/12) is 2(2)1/2/3 .
 
\sin^{-1}\sin5\pi/12 =  \sin^{-1}2(2)1/2/3.
 
i.e.5\pi/12.
 
For, such type of questions we are supposed to know the sin and cos of certain angles or of certain multiples.